OpenStudy (anonymous):

Find all the zeros of each equation. x^3 – x^2 + 5x – 5 = 0 i*sqrt*5 , -i*sqrt*5, 1 i*sqrt*5 , -i*sqrt*5 , -1 *sqrt*5 , -*sqrt*5 , -1 i*sqrt*5 , 1

OpenStudy (anonymous):

x^3+5x-x^2-5=0 x(x^2+5)-1(x^2+5)=0 (x^2+5)(x-1)=0

OpenStudy (anonymous):

x=1, +i sqrt5, -i sqrt 5

OpenStudy (anonymous):

thank you soo much (;

OpenStudy (anonymous):

no problem (;